# Closed Loop Interactive Embodied Reasoning for Robot Manipulation

### v1

@@ III-B Symbolic program generator and action planner
To enable planning of various manipulation tasks we design a set of primitive actions $\mathtt{A}$ that can be easily extended in MuBlE:
@@
  • $\mathtt{squeeze}$ – squeezes the object in the gripper to measure its stiffness.
Non visual states, such as weight or stiffness are encoded in the scene graph after measurements, however the transformer is provided only a binary flag indicating whether the property was measured or not and the sorting reasoning is performed once all the measurements are completed.
III-C Action (render) loop
@@
Scene graph and action planner make part of the action (renderer) loop in Fig. 2. Primitive action controller implements a control that can execute action $\mathtt{A}$ on target object in MuJoCo physics engine or real robot. Physics loop, discussed in more details in Sec. III-D), calculates physics and collects observations $\mathtt{P}$ (e.g. pose of end effector, physical measurement, force in the gripper, etc.). A control signal is generated in every step of the physics loop, until the path is completed. After the execution, a new key frame is captured in real-world setup or rendered in simulation.
Primitive action execution with obstacle avoidance require the positions and orientations $\mathtt{P}$ of the objects in the scene. Every primitive action, including gripper closure, generates a desired trajectory (position and orientation) $\mathtt{P}$ for the end effector. The action of approaching to the grasp position entails a set of pre-coded grasp sequences. We implement grasping of cylindrical objects (depending on their position and size), grasping by the edge, grasping by the handle etc. We use a simplified model as simulated grasping is often not reliable and inconsistent with real grasping. Whenever grasping pads on all gripper fingers are in contact with the same object, it is considered as grasped. Thereafter, the object’s pose is fixed with respect to the end effector. Further, whenever the gripper starts to open, the object is detached from the gripper body with an initial velocity of 0. More details can be found on GitHub page: https://michaal94.github.io/CLIER
III-D Physics loop
The physics are calculated with the use of MuJoCo engine which makes part of MuBlE and is used by CLIER as shown in Fig. 2. With a user defined timestep, the control signal $\mathtt{C}$ is applied to the manipulator, the corresponding forces are calculated and applied to all objects in the scene. Similarly to $\mathtt{robosuite}$, the control signal takes the form of the displacement of the end effector, i.e. position and orientation change, along with binary signal for opening and closing of the gripper. By default, the control signal is translated to desired joint forces by the Operational Space Controller [47], which can be customised. A control signal to the environment can be provided by the user directly or can be obtained from the primitive action controller, which calculates the error of current end effector pose and the planned trajectory.
Every step of applying the control $\mathtt{C}$ affects the physics of the scene and generates a set of observations $\mathtt{P}$, which include: position and orientation of the end effector, joint positions and velocities, status of the gripper closure. Observations of measured non-visual properties, such as stiffness, weight, or elasticity, are also included. Additionally, measurements related to all objects in the scene may be collected, such as positions, orientations, bounding boxes, indication whether the object is currently in the gripper. Any custom sensor can be added to MuBlE as in $\mathtt{robosuite}$.
III-E Reasoning pipeline
CLIER starts by parsing the visual scene to create or update its semantics and geometry scene graph. Simultaneously, a symbolic program in the CLEVR-IEP [25] format is extracted from a natural language instruction. The program is then executed on the scene graph to select the subgoals that have to be achieved on the scene, e.g. measuring the weight of the target when the graph node of the target’s weight is empty. Further, action planning regresses the task to a sequence of primitive actions for the current state of the scene. It reevaluates the next step at every pass of the action (render) loop. The reasoning is completed when either the answer is obtained from the program, or the state of the scene matches the target in task completion modules. The components of the pipeline are presented in Fig. 2 and an example task in Fig. 4.
@@ III-E Reasoning pipeline
Figure 4: An example task from the proposed dataset demonstrating the CLIER pipeline. Blue colour denotes execution of symbolic programs, the output targets from programs are shown in the green, yellow denotes executing sequence of primitive actions after having received a trigger from program (blue). Note that actions in yellow result in updates of the scene graph depicted in the top middle.
IV SHOP-VRB2 Dataset
@@
We introduce SHOP-VRB2 dataset created in MuBlE for training and benchmarking. The dataset includes a set of scenes with instructions to perform various tasks (e.g. Stack metal objects from heaviest to lightest). Tasks are designed to enforce reasoning simultaneously on the visual observations (recognising attributes of objects and their relations) and continuous physical measurements, taking the feedback loop into account. Every example is accompanied with a ground truth sequence of actions for successful execution, along with visual observations and detailed scene graphs. Some examples can be found in Fig. 3 and 4. We include weight measurement as a representative example of estimating non visual object properties through manipulation due to the ease of repeatability for other researchers but we also demonstrate stiffness measurements. Other properties such as roughness are also possible but require another template for training and an implementation of the control. We first discuss the tools we provide in MuBlE that can be used to generate a scene and natural language instructions Sec. IV-A. We then introduce our SHOP-VRB2 benchmark in Sec. IV-B.
IV-A Data generation tools
Scene generator is a tool in MuBlE that can procedurally generate data for training models. We provide a template scene that is compatible with the renderer within the framework. The procedural algorithm takes a set of object models, which can be customised, and randomly places them on the tabletop. Desired properties of the objects such as colour, material, and size can also be randomised. We check for collisions between meshes, unlike previous approaches that check only bounding boxes. Afterwards, a maximal desired level of occlusion is ensured. Finally, we render an image of the scene, generating full ground truth, including segmentation masks for all the objects, the robot, and the tabletop, as well as a depth map. Fig. 3 (Right) presents some example scenes generated with MuBlE. Objects in the scene are described in Sec. IV.
Instruction generator is another tool in MuBlE that given the scene, generates natural language instructions or tasks that require reasoning and interaction. We propose a template-based algorithm where we improve upon CLEVR and SHOP-VRB, which are time-consuming and scale poorly when more adjectives are introduced in the templates. This is due to evaluating all possible combinations of describing words, and validating the descriptions with respect to the given scene. Instead, we implement an efficient rejection mechanism. We first generate all combinations of short descriptions for all scene objects. Short descriptions of objects are more natural than in [48, 13]. Next, we evaluate the descriptions using scene constraints e.g. the target of the instruction has to satisfy the constraint of being pickupable otherwise the description is rejected. The instruction targets are randomly selected from the validated descriptors and used to generate ground truth symbolic program. We automatically produce the ground truth actions by reasoning backwards with predefined pairwise relations e.g. lifting an object requires grasping it, grasping requires approaching to the grasping position, etc.). Fig. 3 presents example instructions generated for the given scenes using templates that are described in Sec. IV.
IV-B Benchmark data
SHOP-VRB2 scenes include $12000$ realistically rendered scenes split to train, validation, and test sets: $10000$:$1000$:$1000$. The scenes contain typical household objects which are easily accessible. Following [13], various instances of objects are included: baking trays, bowls, chopping boards, food containers, glasses, mugs, plates, soda cans, thermoses, and wine glasses. The scenes are generated with $4$ to $5$ objects per scene (their respective distribution is $47.3\%$ and $52.7\%$). We make sure the randomization of the materials (plastic, metal, glass, rubber, wood) is realistic. Out of $54317$ placed objects, a subset was selected randomly, with $1330$ instances of the least common (big chopping board model due to many possible collisions), and $3486$ instances of the most common model (bowl – less prone to collisions).
YCB scenes include 30 simulated benchmarking scenes with 9 YCB-Video [49] objects and 3 randomly generated scenes for each of the benchmarking tasks described below. 9 YCB objects were selected so that the objects share various types of visual/physical attributes (cleanser, mustard, mug, bowl, tomato can, Cheez-it, sugar box, meat can, foam brick).
TABLE II: Instruction templates corresponding to the benchmarking tasks in the proposed dataset.
@@ 2. What is the weight of all $\mathtt{OBJ1}$s?
Benchmarking tasks We assign one instruction to each scene to introduce more diversity in visual observations. We designed 10 classes of benchmarking tasks revolving around moving and stacking objects based on their visual (colour, material, shape), and physical properties (weight). Example templates are presented in Tab. II. The task types are closely related to the instruction templates in Tab. II and involve 1) measuring weight of a single object, 2) measuring weight of multiple objects, 3) picking up based on weight, 4) moving single object 5) moving multiple objects 6) moving based on weight 7) stacking objects 8) stacking objects according to weight 9) stacking 3 objects 10) ordering objects according to their weight. $\mathtt{OBJx}$ refers to a description of an object consisting of a set of visual properties (chosen randomly and validated), e.g. the red object presented in Fig. 1. Note that tasks may refer to either one, specific unique object, or a set of objects sharing a certain property e.g. template 1 and 2 in Tab. II. Further, $\mathtt{TPx}$ specifies a part of the table e.g. the left and the right part. Finally, $\mathtt{WSx}$ refers to weight specifier, i.e. distinction whether the lightest or the heaviest object from the set is the target of the instruction. The length of the instructions ranges between $5$ and $16$ words. The resulting sequences contain between $5$ (measure the weight of the single object) and $46$ (stacking several items according to weight) primitive actions. All instructions are accompanied with symbolic programs in CLEVR-IEP [25] format and task specifications with respect to the scene graph.
Ground truth for visual and physical observations are provided for every scene and its benchmarking instruction. An observation is taken after every primitive action is executed and contains: position and orientation of end effector, gripper status, weight measurement, ground truth scene graph, segmentation masks, and action with its target.
V Benchmarking and results
In this section we provide benchmarking results for CLIER. We show results for two datasets: 1) in simulation for SHOP-VRB2 dataset (Sec. V-A), and 2) on a set of 30 benchmarking scenes in simulation and in real-world with YCB objects (Sec. V-B).
TABLE III: (Left) Success rates for CLIER method on SHOP-VRB2 (VRB, sim) and YCB dataset (sim/real). (Right) Execution outcomes on SHOP-VRB2, incl. successes (bold) and type of failures. See the webpage linked in Sec. I for more details and video illustration.
@@ V Benchmarking and results
Overall $43.9$ $76.7$ $64.4$ Total $100\%$
Benchmarking metrics are the rate of successful task execution (e.g. stacking) and correct question answer (e.g. weight query). Additionally, we provide accuracy split into task types presented in Tab. II. Finally, for the inference tool of CLIER we provide a classification of incorrect attempts, including errors of execution, incorrect scene recognition, loop detection, timeout, inconsistency in object tracking.
@@ V-A SHOP-VRB2 experiments
Results for CLIER are presented in Table III(Left) which shows success rates for tasks. We observe high success rates for the tasks that include manipulation of single objects (weighing or moving one object), and strong decrease in accuracy for multi-object manipulation (stacking more objects, preceding manipulation with weighting). Further, we identify the most common reasons for failure in Table III(Right) which reports the distribution of error codes from the environment. Execution error ($14.4\%$) accounts for failures in execution of primitive actions which may arise from inaccurate scene description (e.g. typically object pose). Scene inconsistency ($12.6\%$) refers to mistakes in tracking objects IDs between frames. Loop detection ($10.8\%$) arises when primitive action chains are repeated (e.g. when approaching object to grasp with misaligned position). Note that this experiment uses simple, ResNet based pose estimation in contrast to more accurate CosyPose [45] that was trained for YCB objects but not for SHOP-VRB2 objects. Finally, we believe that overall accuracy of $43.9\%$ on SHOP-VRB2 indicates that this data forms a challenging benchmark for evaluating visual and interactive reasoning.
V-B Real-world experiments
@@
In this section, we report comparative results for real YCB scenes that mirror 30 YCB scenes introduced in Sec. IV-B.
@@
Experimental Setup includes Franka Emika Panda [50] arm with $7$ degrees of freedom and a $2$-finger parallel gripper. An (extrinsically calibrated) Intel Realsense D455 camera is set to face the robot and capture the front view of the scene (see Fig. 5). To allow CLIER to control the real robot, we developed a thin ZMQ [51] based communication layer that emulates the same interface as for the simulated robot (i.e., motion generation using position-based servoing). Special purpose skills include measuring stiffness via squeezing the object with different forces and measuring deformation or weight via lifting and joint torque differences. Note that the weighing has a relative error of less than ±$\pm 10\%$ in the range $0-200$ g compared to the ground truth weight. The stiffness measurement has a coefficient of variation ranging from $1.6-3.4$, which is highly repeatable. However, we were not able to obtain ground truth data for stiffness.
@@
Sim2Real Our method uses raw RGB rendered images to train the attribute recognition module (c.f. Fig. 2). This is not further tuned between YCB sim and real experiments. The same reasoning model (c.f. Sec. III-E) is evaluated in 30 synthetic and real scenes using the same CosyPose [45] on RGB images. The reasoning modules operate on the scene graph thus are agnostic to visual input.
@@
Figure 5: The real setup with YCB objects (left), corresponding MuJoCo simulation using estimated poses (middle), and RViz visualisation of coloured pointcloud with overlaid grey models detected by CosyPose (right).
@@
Results We mirror experiments with the real scenes in our simulation environment. To evaluate the contribution of different modules we perform testing along with ablations by replacing them with ground truth (GT) data i.e. 1) reasoning success with GT pose and GT attributes 2) pose inference with GT attributes and GT reasoning, 3) pose inference and attribute recognition with GT reasoning, 4) full inference without using GT. The results for all these combinations are in Table IV where the success is the average across 3 runs of each experiment. Note that Seq2seq network in symbolic program generator is not included in the table as it performed with $100\%$ accuracy, therefor the error for case 1) comes from transformer action planner. We observe a very similar performance between simulation and real which validates the usefulness of high quality vision and physics in MuBlE. The performance for individual tasks with YCB scenes in simulation and real world is reported in Tab. III(Left). Sim and Real differences for YCB can be attributed mostly to CosyPose object detection and its robustness to occlusion.
@@
Typical execution errors Our closed loop reasoning approach is able to recover from various execution errors such as dropping or moving an object (see video). Typical errors during the real world execution are caused by uncertainty in pose estimation. This leads to various cases: 1) grasp succeeds because object’s real pose is withing the margin for error. 2) a different grasp is achieved e.g. over a rim instead over diameter of an open cylindrical object. 3) the object slips out when closing the gripper. 4) the gripper collides with the object during the approach and causes a robot error (safety stop reflex). The robot can recover from all except case (4) or when the object is pushed out of the workspace in (3). A varied pose of the object in the gripper (i.e., (1) and (2)) occasionally influences subsequent tasks such as stacking or weighing. Detection of occluded objects is also prone to failures. In particular, occlusion from the gripper when grasping is an issue when updating the scene graph. We use the last estimated object pose and leverage the information of the gripper holding ’something’ to overcome these pose estimation errors.
@@
Joint limits and singularities. Robot kinematic agnostic servoing of the end-effector can lead to degenerate robot configurations. These situations include 1) joint limits, 2) the borders of the workspace, 3) self-collisions, and 4) alignment of several joint axes. To avoid running into joint limits, we start each experiment in a good standard pose and the robot always rotates to a neutral down-facing end-effector orientation before moving via the smallest rotation to the goal rotation. Self collisions are avoided by the Panda controller and are counted as robot errors.
@@
TABLE IV: Success rate for different setups (inference vs. ground truth) in simulation and real world. C - CosyPose, A - attributes recognition, R - reasoning, ✓- module used, ✗- module substituted with ground truth information.

### v2

@@ III-E Primitive actions and physics loop
To enable the planning of various manipulation tasks, we designed a set of primitive actions that can be easily extended in MuBlE:
@@
  • $\mathtt{squeeze}$ – squeezes the object to measure its stiffness.
A specific controller is implemented to execute each primitive action $\mathtt{A}$ on the real robot or in the MuJoCo physics engine within the MuBlE environment (see Sec. IV-A). The action of approaching to the grasp position entails a set of pre-coded grasp sequences. We utilize a simplified grasp simulation within MuBlE – see more details in [14].
Each primitive action is executed within the physics loop. A control signal is generated at every step of the physics loop until the path is completed. It calculates the error for the current end effector pose and the planned trajectory and provides a control signal $\mathtt{C}$ to the environment. Every step of applying the control $\mathtt{C}$ affects the physics of the scene and generates a set of observations $\mathtt{P}$ (e.g. pose of end effector, physical measurement, force in the gripper, etc.). Our method assumes that for each relevant physical property, there exists at least one primitive action that produces observations of this property (e.g. $\mathtt{squeeze}$ for stiffness or $\mathtt{weigh}$ for mass).
III-F Action (render) loop
The action planner reevaluates the next step at every pass (every keyframe) of the action (render) loop that contains also scene graph generation (see Fig. 1). For stability, we let the Action Planner always predict the full sequence of primitive actions until the subgoal completion. After the subgoal is completed (e.g., the weight field is populated in the target node of the graph), the Symbolic Program Executor resumes reasoning. It can either lead to producing the next subgoal (e.g., measuring the weight of another object if the query asks for comparison) or task completion.
@@ III-F Action (render) loop
The Primitive Action Controller implements a control loop that can execute action $\mathtt{A}$ on target object in MuJoCo physics engine or real robot. The physics loop, discussed in Sec. III-E), calculates physics and collects observations $\mathtt{P}$ (e.g. pose of the end effector, physical measurement, force in the gripper, etc.). A control signal is generated in every step of the physics loop, until the path is completed. After the execution, a new key frame is captured in the real-world setup or rendered in simulation.
The reasoning is completed when either the answer is obtained from the program, or the state of the scene matches the target in task completion modules (e.g. Weight of the object is returned or scene graph indicates that objects that were instructed to be stacked are on top of each other).
IV Experimental setup
IV-A MuBlE environment
We implement and evaluate our CLIER method within the MuBlE simulations environment for manipulation tasks [14] (see Fig. 1). It is built on $\mathtt{robosuite}$[43], a framework suitable for creating robotic environments in MuJoCo physics engine. The environment is equipped with high-quality rendering powered by Blender. It is designed for a tabletop scenario with a single robot manipulator and a gripper. It uses the same robots and grippers as $\mathtt{robosuite}$. It provides a set of object models in MuJoCo and their counterparts in Blender, along with templates to create new items.
Physics calculation uses the MuJoCo engine. With a user defined timestep, the control signal $\mathtt{C}$ is applied to the manipulator, the corresponding forces are calculated and applied to all objects in the scene. The simulator provides observations including visual and physical properties of objects and end effector, including measured values for non-visual properties such as weight, stiffness, or elasticity. Any custom sensor can be added to MuBlE as in $\mathtt{robosuite}$.
Data generation tools. The MuBlE environment also provides data generation tools: 1) Scene Generator places selected objects on the scene and sets their properties. It provides the rendered image of the scene (see Fig. 3 (Right)) along with full ground truth data, including segmentation masks as well as a depth map. 2) Instruction Generator given the scene, generates natural language instructions or tasks that require reasoning and interaction. The instruction generator also provides the symbolic program ground truth.
@@ IV-A MuBlE environment
Figure 4: The CLIER pipeline demonstrated on an example task from the SHOP-VRB2 dataset [14]. Blue colour denotes a query and extracted symbolic programs, the output targets from programs are shown in the green, yellow denotes executing the sequence of primitive actions after having received a a subgoal from program (blue). Note that actions in yellow result in updates of the scene graph depicted in the top middle.
IV-B SHOP-VRB2 and YCB-VRB Datasets
@@
We evaluate the proposed CLIER method on SHOP-VRB2 and YCB-VRB benchmarking datasets [14] created in MuBlE [14] for training and benchmarking on long-horizon manipulation tasks. The datasets include a set of scenes with instructions to perform various tasks (e.g. Stack metal objects from heaviest to lightest). Tasks are designed to enforce reasoning simultaneously on the visual observations (recognising attributes of objects and their relations) and continuous physical measurements (e.g. measuring invisible properties), taking the feedback loop into account. Every example is accompanied with a ground truth sequence of actions in CLEVR-IEP [16] format, along with visual observations and detailed scene graphs, including physical properties (see examples in Fig. 3 and 4).
IV-B1 Benchmarking tasks
There are 10 classes of benchmarking tasks (see Tab. I; more details can be found at [14]). $\mathtt{OBJx}$ refers to a description of an object including visual properties, e.g. the red object presented in Fig. 3 (Right). $\mathtt{TPx}$ specifies left/right part of the table and $\mathtt{WSx}$ is a weight specifier (lightest/heaviest). Instructions have $5$ to $16$ words. The resulting sequences contain between $5$ (measure the weight of a single object) and $46$ (stacking several items according to weight) primitive actions.
IV-B2 SHOP-VRB2 dataset
SHOP-VRB2 dataset [14] extends SHOP-VRB [38] dataset by introducing non-visible object properties and instructions that require closed-loop reasoning. SHOP-VRB2 scenes include $12000$ realistically rendered scenes with typical household objects (similarly to [38]). The scenes are generated as described in MuBlE [14]. Scene generator with $4$ to $5$ objects per scene. SHOP-VRB2 instructions correspond to the tasks described in Sec. IV-B2 and are generated using MuBlE Instruction Generator tool (see Sec. IV-A). We assign one instruction to each scene. The instructions are randomly generated, however, they keep an equal distribution among the tasks from Table I and preserve training/test split as suggested in [14].
TABLE I: Instruction templates and benchmarking tasks in SHOP-VRB2.
@@ 2. What is the weight of all $\mathtt{OBJ1}$s?
IV-B3 YCB-VRB dataset
YCB-VRB scenes include 30 simulated benchmarking scenes with 4-5 objects randomly selected from the set of 9 YCB-Video [44] objects sharing various visual/physical attributes. YCB-VRB Instructions are generated similarly to SHOP-VRB2. We generate 3 instructions per every template from Table I.
IV-B4 Non-visual properties
Although the SHOP-VRB2 and YCB-VRB datasets contain only weight measurement as a representative example of estimating non-visual object properties through manipulation due to the ease of repeatability for other researchers, we also demonstrate performance of our method on tasks requiring stiffness measurements (see the accompanying video [45]) to show both the ease of extending the benchmark to other physical properties as well as the applicability of the proposed CLIER method to various types of embodied actions. Including such a novel property only requires adding another instruction template for training and implementation of the control action to measure the given physical property.
V Results
In this section we provide results for CLIER: 1) in simulation for SHOP-VRB2 dataset (Sec. V-A), and 2) on a set of 30 benchmarking scenes in simulation and in the real world with YCB objects (Sec. V-B).
TABLE II: (Left) Success rates for CLIER method on SHOP-VRB2 (SHOP, sim) and YCB dataset (YCB sim/real). (Right) Execution outcomes on SHOP-VRB2, incl. successes (bold) and types of failures. See the webpage linked in Sec. I for more details and video illustration.
@@ V Results
Overall $43.9$ $76.7$ $64.4$
Evaluation metrics are the rate of successful task execution (e.g. stacking) and percentage of correctly answered questions (e.g. weight query). We also provide accuracy split into task types presented in Tab. I. Finally, for the inference tool of CLIER we provide a classification of incorrect attempts, including errors of execution, incorrect scene recognition, loop detection, timeout, and inconsistency in object tracking.
@@ V-A SHOP-VRB2 experiments
Results for CLIER are presented in Table II (Left) which shows success rates for various tasks. We observe high success rates for the tasks that include manipulation of single objects (weighing or moving one object) and a strong decrease in accuracy for multi-object manipulation (stacking more objects, with preceding manipulation for weighting). Further, we identify the most common reasons for failure in Table II (Right), which reports the distribution of error codes from the environment. Execution error ($14.4\%$) accounts for failures in execution of primitive actions which may arise from inaccurate scene description (e.g. typically object pose). Scene inconsistency ($12.6\%$) refers to mistakes in tracking objects IDs between frames. Loop detection ($10.8\%$) arises when primitive action chains are repeated (e.g. when approaching object to grasp with misaligned position). Note that this experiment uses simple, ResNet based pose estimation in contrast to more accurate CosyPose [41] that was trained for YCB objects but not for SHOP-VRB2 objects. Finally, we believe that the overall accuracy of $43.9\%$ on SHOP-VRB2 indicates that this data forms a challenging benchmark for evaluating visual and interactive reasoning.
V-B YCB-VRB and real-world experiments
@@
In this section, we report comparative results for real YCB scenes that mirror 30 YCB scenes introduced in Sec. IV-B2.
@@
Experimental Setup includes the Franka Emika Panda [46] arm with $7$ degrees of freedom and a $2$-finger parallel gripper. An (extrinsically calibrated) Intel Realsense D455 camera is set to face the robot and capture the front view of the scene (see Fig. 5). To allow CLIER to control the real robot, we developed a thin ZMQ [47] based communication layer that emulates the same interface as for the simulated robot (i.e., motion generation using position-based servoing). Special purpose skills include measuring stiffness via squeezing the object with different forces and measuring deformation or weight via lifting and joint torque differences. Note that the weighing has a relative error of less than ±$\pm 10\%$ in the range $0-200$ g compared to the ground truth weight. The stiffness measurement has a coefficient of variation ranging from $1.6-3.4$, which is highly repeatable. However, we were not able to obtain ground truth data for stiffness.
@@
Sim2Real. Our method uses raw RGB rendered images to train the attribute recognition module (see Fig. 1). This is not further tuned between the YCB sim and real experiments. The same reasoning model (see Sec. III) is evaluated in 30 synthetic and real scenes using the same CosyPose [41] on RGB images. The reasoning modules operate on the scene graph and thus are agnostic to visual input.
@@
Figure 5: The real setup with YCB objects (left), corresponding MuJoCo simulation using estimated poses (middle), and RViz visualisation of colored pointcloud with overlaid gray models detected by CosyPose (right).
@@
Results. We mirror experiments with real scenes in our simulation environment. To evaluate the contribution of different modules we perform testing along with ablations by replacing them with ground truth (GT) data, i.e. 1) reasoning success with GT pose and GT attributes 2) pose inference with GT attributes and GT reasoning, 3) pose inference and attribute recognition with GT reasoning, 4) full inference without using GT. The results for all these combinations are in Table III, where the success is the average across 3 runs of each experiment. Note that the Seq2seq network in the symbolic program generator is not included in the table as it performed with $100\%$ accuracy, therefore the error for case 1) comes from the transformer action planner. We observe a very similar performance between simulation and real world which validates the usefulness of high-quality vision and physics in MuBlE. The performance for individual tasks with YCB scenes in simulation and the real world is reported in Tab. II(Left). Sim and Real differences for YCB can be attributed mostly to CosyPose object detection and its robustness to occlusion.
@@
Typical execution errors. Our closed loop reasoning approach is able to recover from various execution errors such as dropping or moving an object (see video [45]). Typical errors during the real world execution are caused by uncertainty in pose estimation. This leads to various cases: 1) grasp succeeds because object’s real pose is withing the margin for error. 2) a different grasp is achieved e.g. over a rim instead over diameter of an open cylindrical object. 3) the object slips out when closing the gripper. 4) the gripper collides with the object during the approach and causes a robot error (safety stop reflex). The robot can recover from all except case (4) or when the object is pushed out of the workspace in (3). A varied pose of the object in the gripper (i.e., (1) and (2)) occasionally influences subsequent tasks such as stacking or weighing. The detection of occluded objects is also prone to failures. In particular, occlusion from the gripper when grasping is an issue when updating the scene graph. We use the last estimated object pose and leverage the information of the gripper holding ‘something’ to overcome these pose estimation errors.
@@
Joint limits and singularities. Cartesian control of the robot end-effector without taking the joint configuration into account can lead to degenerate configurations. These include 1) joint limits, 2) borders of the workspace, 3) self-collisions, and 4) alignment of several joint axes. To avoid running into joint limits, we start each experiment in a good standard pose, and the robot always rotates to a neutral down-facing end-effector orientation before moving via the smallest rotation to the goal rotation. Self collisions are avoided by the Panda controller and are counted as robot errors.
@@
TABLE III: Success rate for different setups (inference) in simulation and real world. C - CosyPose, A - attributes recognition, R - reasoning, ✓- module used, ✗- module substituted with ground truth information.